1. If the electric field in an EM wave in free space has a peak magnitude of 8.7×10−5V/m, what is the peak magnitude of the magnetic field strength?

B0 = _________ T

2. A 1500-nF capacitor with circular parallel plates2.0 cm in diameter is accumulating charge at the rate of 39.0 mC/s at some instant in time.

a) What will be the induced magnetic field strength 12.5 cm radially outward from the center of the plates?

B1 = _______ T

b) What will be the value of the field strength after the capacitor is fully charged?

B2 = _______ T

3. In an EM wave traveling west, the B field oscillates up and down vertically and has a frequency of 95.0 kHz and an rms strength of7.85×10−9T. Assume that the wave travels in free space.

a) What is the frequency of the electric field?

b) What is the rms strength of the electric field?

c) What is the direction of its oscillation?

()The electric field oscillates along the horizontal north-south line.
()The electric field oscillates along the horizontal west-east line.
()The electric field oscillates vertically.
()None of the above.

4. A light-year is a measure of distance (not time). How many meters does light travel in a year? One year is 3.156×107s.

d= ____________

5. An EM wave in free space has frequency 6.99×1014Hz.

a) What is its wavelength?

b) How would we classify it?

()radio
() infrared
() visible
() ultraviolet
() X-rays
() gamma

6. A widely used "short-wave" radio broadcast band is referred to as the 75-m band. What is the frequency of a 75-m radio signal?

f= _________

7. Calculate the displacement current ID between the square plates, 7.6 cm on a side, of a capacitor if the electric field is changing at a rate of 2.0×106 V/m⋅s.

ID = __________

8. How long does it take light to reach us from the Sun, 1.50 ×108km away?

t = __________ min

9. An EM wave in free space has a wavelength of630 nm .

a) What is its frequency?

f= _____________

b) How would we classify it?

() infrared
() ultraviolet
() X-rays
() gamma
() visible
() radio

10.

a) What is the wavelength of a 27.25×109Hz radar signal in free space? The speed of light is 2.9979×108m/s.

λ = ________________

b) What is the frequency of an X-ray with wavelength 0.12 nm ? Assume that the wave travels in free space.

f = __________________
Answer #1

1) Given,

Em = 8.7×10-5 V/m

We need to find the magnitude of magnetic field.

We know that,

c = Em/Bm => Bm = Em/c

Bm = 8.7 x 10-5 V/m/ 3 x 108 m/s

Bm = 2.9 x 10-13 T

Hence, Bm = 2.9 x 10-13 T.

5)Given

Frequency = 6.99 x 1014 Hz

a) Let \lambda be the wavelength.

We know that the wavelength, speed and frequency of waves are related as

v = f \lambda => \lambda = v/f = c/f

\lambda = 3 x 108 m/s /  6.99 x 1014 Hz = 0.429 x 10-6 m

Hence, \lambda =  0.429 x 10-6 m = 429 x 10-9 m = 429 nm

b)this falls under visible spectrum.

[PS:- Humbly requesting you to put the other questions seperately].